3. The Peterson’s paid for many years on a $60,000 mortgage with a term of 30 years and a 4.5% interest rate. Prepare an amortization schedule for the first 2 months of their mortgage.
Answer by Theo(13342)

the present value of the mortgage is 60,000.

the term is 30 years and the interest rate is 4.5% per year.

the payments are monthly, therefore it can be assumed that the interest rate is compounded monthly since the prevailing standard is that the number of discounting periods per year is the same as the number of payments per year.

use a financial calculator to find that the monthly payment will be 304.l0111859 when payments are made at the end of each month.

the following excel printout shows the first few months of making these payments and the remaining balance at the end of each month that results.

$$$

this printout shows that:

at the end of the first month (tp 1), the remaining balance is $59,920.99.

at the end of the second month (tp 2), the remaining balance is $59,841.68.

using a financial calculator, the monthly payment was determined to be equal to approximately $304.01 per month.

the monthly interest rate was determined to be 4.5/1200 = .00375 per month.

the calculations for the end of the first month were:

60,000 * 1.00375 - 304.01 = 59,920.99.

the calculations for the end of the second month were:

59,920.99 * 1.00375 - 304.01 = 59,841.68.

the excel printout confirms these calculations to correct, as it should, since it used the same formula.

the last few months of the mortgage calculations are shown below:

$$$

as you can see, the remaining balance is 0 at the end of the 360th month.

this confirms the calculation of the monthly payment was correct.

i used a financial calculator to find the monthly payment.

using this calculator, i entered the following:

present value = -60,000
interest rate per month = 4.5 / 1200 = .375%
number of months = 30 * 12 = 360
payments are made at the end of the time period.
payment per month = 0
future value = 0

i then instructed the calculator to find the monthly payment per month.

a similar calculation using an online financial calculator is shown below:

$$$

this calculator can be found at https://arachnoid.com/finance/